# arXiv:2603.00415v1[eess.SY] 28 Feb 2026

### v1

@@ Abstract
Artificial intelligence (AI) is driving a rapid expansion of data centers (DCs). These facilities consume large amounts of electricity and introduce new challenges for power systems. AI workloads cause rapid power changes and high peak demand. These behaviors are different from traditional data centers (TDCs) and can affect grid stability and reliability. This paper reviews how energy storage systems (ESSs) can help integrate AI data DCs with the electric grid. We examine storage solutions at multiple levels, including grid-scale batteries, UPS systems, rack-level storage, and chip-level buffering. Each layer operates at a different time scale and serves a different purpose. Grid-interactive UPS (GiUPS) systems can respond quickly to disturbances and assist with frequency regulation or voltage ride through. Large battery energy storage systems (BESSs) can smooth power demand, support renewable on-site generation, and provide grid services. Rack-level and server-level storage help manage fast power fluctuations close to computing hardware. We also discuss other technologies such as fuel cells (FCs) and thermal energy storage (TE) that can support co-generation and reduce emissions. In addition, second-life battery energy storage (SLBESS) are reviewed as a lower-cost option for large installations whether supporting UPS battery or as a backup generation. The paper compares the benefits, challenges, and coordination requirements of these solutions. Overall, the study provides a structured view of how energy storage can improve reliability, flexibility, and sustainability when connecting future AI data centers to the power grid.
Index Terms:
AI data center, battery backup unit (BBU), battery energy storage systems (BESS), fuel cell (FC), GPU, second-life battery energy storage systems (SLBESS), thermal energy storage (TES), UPS
@@ I Introduction
Artificial Intelligence (AI) has fundamentally altered society’s relationship with technology; spanning applications from simple search tasks to advanced scientific discovery. Large language models (LLMs) have been at the forefront of these high-complexity AI systems, with models such as ChatGPT and Claude, increasingly embedded in everyday use across the United States. These large-scale models are trained on trillions of text tokens collected from online sources and are deployed through web-based platforms that can attract millions over short periods of time. The surge in demand has prompted AI industry stakeholders to consider contemporary data center (DC) designs capable of supporting the high-computational tasks required for training, inference, and fine-tuning of increasingly more complex models. Energy requirements for AI queries are now nearly an order of magnitude greater than those of Google search, while AI training workloads have been observed to double approximately every 3.4 months [15]. Collectively, these trends highlight the necessity for rethinking data center architectures to sustainably support the rapid scaling of AI workloads.
@@
In contrast, AI DCs are driven by compute-intensive workloads, the widespread deployment of power-hungry GPUs, and the use of high-efficiency cooling technologies, all of which significantly increase power requirements. Thus, large-scale AI training facilities are increasingly located in remote or semi-rural regions where land, power, and cooling resources are more accessible and cost-effective. Rack-level power demand in AI DCs can exceed 100 kW, while aggregate facility consumption may scale to hundreds of megawatts or even gigawatt levels [57]. As a result, large-scale AI training facilities often require direct interconnection with transmission networks and face more strict reliability, siting, and environmental constraints than traditional facilities [121]. Inference-oriented facilities, while potentially smaller in individual footprint, may connect at either distribution or transmission level depending on aggregate scale and siting.
@@
Beyond the differences in scale and infrastructure, AI DCs exhibit two distinct load profiles (training and inference) that differ fundamentally from those of TDCs. AI training workloads are characterized by rapid power fluctuations arising from idle periods, peak utilization events, and checkpointing operations, with significant load variations occurring on sub-second timescales. In addition, training jobs often produce high-frequency, jitter-like power spikes. In contrast, inference workloads generally lack such rapid transients and exhibit comparatively smoother demand profiles. TDC loads, by comparison, behave more like conventional grid loads and are typically predictable using historical data and standard forecasting methods.
@@
AI DCs are connected to the power grid through high-density power converters, which lead to higher power distortion at the utility level. Besides, jittery spikes in training jobs may cause sub-frequency mechanical oscillations in synchronous generators, which increase mechanical stress and accelerate corrosion. As a result, harmonic compensation and load smoothing become a critical requirement for AI DC operation, both to improve load predictability and to mitigate adverse power-quality impacts on the power system.
@@
Large-scale AI DCs enforce the power grid to accommodate higher reserve margins and proactive resource planning. This is because forecasting their load dynamics is more challenging than for traditional loads, and the uncertainty associated with these types of loads is significantly higher. Therefore, additional reserve capacity is necessary to maintain normal grid operation and to support grid expansion when integrating AI DCs.
@@
Connecting or disconnecting large-scale AI DCs may jeopardize system stability due to high-range frequency fluctuations. These facilities are designed to rapidly transfer to backup power in response to even minor grid disturbances, which can trigger large, sudden load drops at the transmission level. Such disturbances may activate load-frequency relays or enforce grid operators to implement load-shedding scenarios. A recent transmission fault in Virginia triggered an abrupt 1,500 MW reduction in system load as multiple DCs transitioned to backup power, leading to significant frequency and voltage deviations that required operator intervention [80]. Events such as this illustrate the limitations of conventional grid planning and operational strategies when accommodating large AI DC loads.
@@
Considering the above-mentioned challenges, we investigate on the role and impact of energy storage systems (ESSs) in enabling the smooth integration of future large-scale AI DCs into the power system. To the best of our knowledge, there is no comprehensive document specifically focusing on the potential support roles of ESSs in the reliable integration of AI DCs with the power grid. In this paper, we address this gap and investigate the role of ESS in AI DCs. ESSs can act not only as backup units with limited uptime but also support cooling through thermal storage units. They can be deployed as battery backup units (BBUs) inside IT racks, GPU-level batteries for chip-level load smoothing, and battery energy storage systems (BESSs) for DC-level load smoothing or co-generation support. ESSs within UPS systems can also interact with the utility grid to provide services such as frequency regulation and peak shaving. Fuel cells (FCs) are also have the potential to replace diesel back-up generation with technology advancements.
@@
In contrast to [16, 46], this paper is highly focused on the role of ESSs in enabling reliable grid integration of AI DCs, rather than only characterizing the grid impacts of these large loads. Moreover, our work does not focus on ESS chemistry technologies [108, 114], but instead attempts to reveal design architectures for the optimal application of ESSs in AI DCs grid integration. Therefore, this study serves as an appropriate source for both academic and industry partners seeking the required knowledge to explore new avenues for the role of ESSs in future AI DCs.
The remainder of the paper is organized as follows. Section II provides an overview of the AI DC architecture, where IT equipment, power delivery systems, cooling systems, and miscellaneous units are examined, along with the required upgrades compared to TDCs. Section III investigates the role and operation of UPS systems, with particular emphasis on the grid-interactive UPS (GiUPS) systems, which has recently been studied as a key component of future AI DCs. Section IV discusses the integration of BESSs with AI DCs and their role in power smoothing and reliable grid integration, with a special focus on large-scale units. Sections V and VI address FCs and thermal energy storage (TES) systems, respectively, as non-battery-based energy storage solutions, highlighting their contributions to low-carbon on-site generation and more efficient cooling systems while maximizing the utilization of renewable energy sources (RESs). The role of rack-level and server-level ESSs is also examined as primary components for mitigating chip- and server-level power fluctuations, and their coordination requirements are emphasized in section VII. Section VIII investigates the role of second-life BESSs (SLBESSs) as cost-effective ESSs within the AI-DC architecture, and presents evidence demonstrating their capability to support centralized BESSs in AI DCs. Finally, Section X discusses the challenges associated with integrating multi-layer ESSs and outlines potential future research directions.
@@ II AI DC Power Architecture
The AI DC infrastructure comprises different sections, including IT equipment, power distribution networks, cooling systems, and miscellaneous units such as lighting and security. In this section, we provide a brief introduction to the AI DC components and the ESS units in its structure. The overall structure of the AI DC with all ESSs is provided in Fig. I.
@@ II-C Cooling Systems
TABLE I: Overview of GiUPS Operating Modes
@@ III-B1 Operational Modes of GiUPS
GiUPS can operate in several modes depending on the grid conditions and external control signals, including standard operation, discharge modes (full and partial disconnection), recharge mode, and energy export (bi-directional) mode [31, 35]. When a frequency variation is detected by an external controller, the GiUPS adjusts its operation to provide both positive and negative regulation by charging or discharging its batteries within operational limits. Table I, shows an overview of GiUPS operation modes.
@@ III-C BESS Sizing in GiUPS Systems
The sizing of the BESS in GiUPS applications must simultaneously satisfy backup reliability requirements and grid-service performance objectives [46, 108]. GiUPS participation in services such as FFR, primary regulation, and peak shaving introduces additional sizing considerations. Frequency regulation is primarily power-intensive rather than energy-intensive, requiring high ramp rates and short-duration injections. Consequently, the inverter power rating may approach the full UPS capacity even if incremental energy capacity beyond backup requirements remains moderate. Cycle life and cumulative energy throughput are critical determinants of economic sizing. Grid-service participation may introduce hundreds of shallow cycles annually [104, 58, 145, 133]. Battery degradation correlates strongly with equivalent full cycles, temperature, and average state of charge [59, 97, 74, 33]. Moderate oversizing of the BESS can reduce effective depth of discharge (DOD) per cycle, extending service life and lowering long-term replacement cost. However, excessive oversizing increases capital expenditure, footprint, and thermal management complexity without proportionate revenue gains.
An optimal sizing methodology therefore integrates (i) backup energy constraints, (ii) committed grid-service power capacity, and (iii) degradation-aware lifecycle cost modeling. Optimization frameworks typically minimize total cost of ownership while ensuring reliability margins and projected ancillary service revenues. For AI DCs with highly dynamic loads, an additional allocation of BESS capacity may be reserved for short-term load smoothing to mitigate upstream ramp rates and transformer stress. This buffering function must be explicitly included in sizing calculations to prevent erosion of guaranteed backup autonomy. This further highlights the role of long-duration ESSs (LDESSs) in modern UPS systems, which are currently under investigation for grid use cases.
@@ IV-A BESS Control Modes
IV-A1 GFM Control Model
In grid-forming operation, the BESS behaves as a controlled voltage source and establishes its own internal frequency and phase without relying on an external phase-locked loop (PLL) [66]. The controller emulates the swing dynamics of a synchronous generator by incorporating virtual inertia, damping, and active power–frequency droop characteristics.
The inverter angle dynamics are given by
$\frac{d\delta_{\mathrm{vsm}}}{dt}=\omega_{\mathrm{vsm}},$ (1)
while the internal frequency dynamics are described by
$\small 2H_{\mathrm{vsm}}\frac{d\omega_{\mathrm{vsm}}}{dt}=P_{\mathrm{ref}}-P-\left(\frac{1}{m_{p}}+D_{1}+\frac{sD_{2}}{s+\omega_{D}}\right)\left(\omega_{\mathrm{vsm}}-\omega_{0}\right).$ (2)
Here, $H_{\mathrm{vsm}}$ is the virtual inertia constant, $m_{p}$ is the active power–frequency droop coefficient, $D_{1}$ represents steady-state damping, $D_{2}$ is a transient damping term shaped by a washout filter with corner frequency $\omega_{D}$, $\omega_{\mathrm{vsm}}$ is the inverter angular frequency, $\omega_{0}$ is the nominal angular frequency, $P_{\mathrm{ref}}$ is the active power reference, and $P$ is the measured inverter output power. This formulation enables autonomous frequency regulation and inertial response, making GFM BESSs suitable for islanded operation [82].
IV-A2 GFL Control Model
In contrast, GFL BESS operate as controlled current sources that synchronize to an externally imposed grid voltage using a PLL [10]. The inverter frequency and phase are therefore dictated by the grid rather than internally generated.
A conventional synchronous-reference-frame PLL can be expressed as
$\omega_{\mathrm{pll}}=\omega_{0}+K_{p}v_{q}+K_{i}\int v_{q}\,dt,$ (3)
$\frac{d\delta_{\mathrm{pll}}}{dt}=\omega_{\mathrm{pll}},$ (4)
where $v_{q}$ is the quadrature-axis component of the measured grid voltage in the PLL reference frame, $K_{p}$ and $K_{i}$ are the proportional and integral gains of the PLL, and $\delta_{\mathrm{pll}}$ and $\omega_{\mathrm{pll}}$ denote the estimated grid angle and frequency, respectively.
Because GFL inverters depend on a strong external voltage reference, their performance degrades under weak-grid conditions, motivating the adoption of GFM control strategies in inverter-dominated systems [82].
The BESS terminal-voltage dynamics $E$ are modeled as
$k_{iv}\frac{dE}{dt}=V_{\mathrm{ref}}+m_{q}\left(Q_{\mathrm{ref}}-Q\right),$ (5)
where $k_{iv}$ is the integral gain, $V_{\mathrm{ref}}$ is the voltage reference, $m_{q}$ is the QQVV droop coefficient, $Q_{\mathrm{ref}}$ is the reactive power reference, and $Q$ is the measured reactive power. This formulation regulates the inverter voltage magnitude through a reactive-power droop relationship.
TABLE II: Summary of power-smoothing architectures for AI data centers: benefits and challenges [107]
@@ IV-B2 Power Smoothing
As shown in Fig. 2, the GFM BESS effectively smooths the power fluctuations of the AI DCs resulting from training and inference jobs. An alternative approach is a hybrid E-STATCOM and BESS configuration, which combines the fast, high-speed charge/discharge response of supercapacitors with the higher energy capacity of BESS to achieve effective smoothing and sustained load support. This hybrid architecture can enhance resilience by improving demand flexibility, reducing flicker, and enabling broader grid-service capability. The primary trade-off is increased cost and footprint compared with standalone solutions. In addition, standalone supercapacitor-based E-STATCOM solutions can mitigate high-frequency AI load fluctuations and improve power quality for sensitive IT equipment by reducing flicker, although they provide limited long-duration energy support [107]. Table II compares the benefits and challenges of representative BESS-based and hybrid strategies for enabling grid support from AI DCs.

### v2

# Grid Integration of AI Data Centers: A Critical Review of Energy Storage Solutions

@@ Abstract
Artificial intelligence (AI) is driving unprecedented growth in data center (DC) scale and power demand. AI workloads impose highly dynamic, difficult-to-forecast power profiles on the utility grid, creating reliability and stability challenges that conventional DC architectures are not designed to address. This paper provides a critical review of energy storage systems (ESSs) as the key enabling technology for reliable grid integration of AI DCs. We organize the review around a four-layer hierarchical taxonomy, namely chip-level buffering, rack/server-level ESSs, facility-level uninterruptible power supply (UPS) systems, and grid-scale battery energy storage systems (BESSs), supplemented by non-battery technologies including fuel cells (FCs) and thermal energy storage (TES). Each layer is analyzed with respect to response timescale, power and energy ratings, operational role, integration challenges, and coordination requirements. Key findings include: (i) AI DC load profiles differ fundamentally from traditional loads in their sub-second variability, making conventional ESS dispatch strategies insufficient; (ii) hierarchical, coordinated ESS deployment across all layers is necessary for effective load smoothing and grid support; and (iii) significant gaps remain in simulation tools, degradation modeling, load forecasting, and optimal multi-layer sizing. This review identifies open research challenges and future directions at the intersection of AI computing infrastructure and power system integration.
Abbreviations and Symbols
@@
RoCoF Rate of Change of Frequency
RUL Remaining Useful Life
SEI Solid Electrolyte Interphase
@@
Key Symbols
$C_{ch,t}$ Electricity price during charging
at time step $t$ (\$/kWh)
$C_{deg}$ Degradation cost coefficient (\$/cycle)
$C_{dis,t}$ Electricity price during discharging
at time step $t$ (\$/kWh)
$P_{ch}$ Charging power (kW)
$P_{dis}$ Discharging power (kW)
$t$ Discrete time step (hours)
$\Delta f$ Frequency deviation (Hz)
@@ I Introduction
Artificial Intelligence (AI) has fundamentally altered society’s relationship with technology, spanning applications from simple search tasks to advanced scientific discovery. Large language models (LLMs) have been at the forefront of these high-complexity AI systems, with models such as ChatGPT and Claude becoming increasingly embedded in everyday use across the United States. These large-scale models are trained on trillions of text tokens collected from online sources and are deployed through web-based platforms that can attract millions of users over short periods of time. This surge in demand has prompted AI industry stakeholders to reconsider contemporary data center (DC) designs so that they can support the intensive computational requirements of training, inference, and fine-tuning for increasingly complex models. Energy requirements for AI queries are now nearly an order of magnitude greater than those of Google search, while AI training workloads have been observed to double approximately every 3.4 months [14]. Collectively, these trends highlight the need to rethink data center architectures to sustainably support the rapid scaling of AI workloads.
@@
In contrast, AI DCs are driven by compute-intensive workloads, the widespread deployment of power-hungry GPUs, and the use of high-efficiency cooling technologies, all of which substantially increase power requirements. As a result, large-scale AI training facilities are increasingly located in remote or semi-rural regions where land, power, and cooling resources are more accessible and cost-effective. Rack-level power demand in AI DCs can exceed 100 kW, while aggregate facility consumption may scale to hundreds of megawatts or even gigawatt levels [56]. Consequently, large-scale AI training facilities often require direct interconnection with transmission networks and face stricter reliability, siting, and environmental constraints than traditional facilities [119]. Inference-oriented facilities, while potentially smaller in individual footprint, may connect at either the distribution or transmission level depending on aggregate scale and siting.
Figure 1: Overall architecture of an AI data center with hierarchical energy storage systems. The figure illustrates the power delivery path from the utility grid through the medium-voltage transformer, UPS/GiUPS, power distribution units, and rack busbars to the IT equipment (GPUs/TPUs).
@@
Beyond differences in scale and infrastructure, AI DCs also exhibit two distinct load profiles, training and inference, that differ fundamentally from those of TDCs. AI training workloads are characterized by rapid power fluctuations arising from idle periods, peak utilization events, and checkpointing operations, with significant load variations occurring on sub-second timescales [71, 128]. In addition, training jobs often produce high-frequency, jitter-like power spikes. In contrast, inference workloads generally lack such rapid transients and exhibit comparatively smoother demand profiles. TDC loads, by comparison, behave more like conventional grid loads and are typically predictable using historical data and standard forecasting methods.
@@
These distinctive load characteristics have important implications for grid operation. AI DCs are connected to the power grid through high-density power converters, which can increase power distortion at the utility level [68, 140]. In addition, jittery spikes in training jobs may cause sub-frequency mechanical oscillations in synchronous generators, thereby increasing mechanical stress and accelerating corrosion. As a result, harmonic compensation and load smoothing become critical requirements for AI DC operation, both to improve load predictability and to mitigate adverse power-quality impacts on the power system.
@@
Large-scale AI DCs also require the power grid to accommodate higher reserve margins and more proactive resource planning [15]. This is because forecasting their load dynamics is more challenging than for traditional loads, and the uncertainty associated with these facilities is significantly higher. Therefore, additional reserve capacity is needed both to maintain normal grid operation and to support grid expansion when integrating AI DCs.
@@
Moreover, connecting or disconnecting large-scale AI DCs may jeopardize system stability because of high-range frequency fluctuations. These facilities are designed to rapidly transfer to backup power in response to even minor grid disturbances, which can trigger large and sudden load drops at the transmission level. Such disturbances may activate load-frequency relays or force grid operators to implement load-shedding actions. A recent transmission fault in Virginia triggered an abrupt 1,500 MW reduction in system load as multiple DCs transitioned to backup power, leading to significant frequency and voltage deviations that required operator intervention [79]. Events such as this illustrate the limitations of conventional grid planning and operational strategies when accommodating large AI DC loads.
@@
Considering these challenges, this paper investigates the role and impact of energy storage systems (ESSs) in enabling the smooth integration of future large-scale AI DCs into the power system. To the best of our knowledge, there is no comprehensive document specifically focused on the potential support roles of ESSs in the reliable integration of AI DCs with the power grid. In this context, ESSs can serve not only as backup units with limited uptime, but also as enablers of cooling support through thermal storage units. They can be deployed as battery backup units (BBUs) inside IT racks, GPU-level batteries for chip-level load smoothing, and battery energy storage systems (BESSs) for data-center-level load smoothing or co-generation support. ESSs embedded within UPS systems can also interact with the utility grid to provide services such as frequency regulation and peak shaving. In addition, fuel cells (FCs) have the potential to replace diesel backup generation as the technology matures.
To provide a structured perspective on the diverse storage technologies reviewed in this paper, Table I summarizes the hierarchical ESS classification framework organized by storage layer, typical response timescale, power and energy ratings, primary application role, and the section in which each technology is discussed. This classification is intended to guide readers in comparing technologies across layers and to serve as a roadmap for the remainder of the paper. A more detailed cross-technology comparison, including cost and suitability metrics, is provided in Table V.
@@
In contrast to [15, 45], this paper focuses specifically on the role of ESSs in enabling reliable grid integration of AI DCs, rather than only characterizing the grid impacts of these large loads. Moreover, our work does not focus on ESS chemistry technologies [106, 112], but instead aims to reveal design architectures for the optimal application of ESSs in AI DC grid integration. Therefore, this study serves as an appropriate source for both academic and industry partners seeking the knowledge required to explore new directions for the role of ESSs in future AI DCs. The main contributions of this review paper are summarized as follows:
  • Hierarchical Taxonomy: We propose a four-layer classification framework organizing ESSs in AI DCs by storage level, response timescale, and application role: (i) chip-level power buffers (sub-millisecond to milliseconds), (ii) rack/server-level BBUs and supercapacitors (milliseconds to seconds), (iii) facility-level GiUPS systems (seconds to minutes), and (iv) grid-scale BESSs (minutes to hours). This taxonomy provides a structured basis for comparing and co-designing ESSs across layers.
  • Grid-Interactive UPS Analysis: We provide a detailed review of GiUPS operating modes, control architectures, grid-service capabilities (FFR, frequency regulation, VRT, peak shaving), and BESS sizing methodologies, an area that has received limited systematic treatment in prior literature.
  • Comparative Discussion of ESS Technologies: We critically compare battery-based ESSs, fuel cells, thermal storage, and second-life batteries with respect to their response time, power and energy ratings, application level, advantages, limitations, cost considerations, and suitability for AI DC integration (see Table V).
  • Non-Battery ESS Review: We review the roles of FCs and TES, which are often neglected in AI DC literature, as complementary technologies supporting low-carbon on-site generation and advanced cooling, respectively.
  • Challenges and Research Gaps: We systematically identify open challenges in simulation and validation, GPU scheduling, degradation modeling, load forecasting, hierarchical coordination, and long-duration storage, and propose specific future research directions.
@@ II AI DC Power Architecture
The AI DC infrastructure comprises different sections, including IT equipment, power distribution networks, cooling systems, and miscellaneous units such as lighting and security. In this section, we provide a brief introduction to the AI DC components and the ESS units in its structure. The overall structure of the AI DC with all ESSs is provided in Fig. 1.
TABLE I: Hierarchical ESS Classification Framework for AI Data Centers
Layer Technology Response Timescale Typical Rating Primary Role Section
Chip-level On-chip capacitors, HBM buffers Sub-ms to ms W–kW GPU power spike smoothing VII
Rack/Server-level BBUs, supercapacitors, rack BESS ms to seconds kW–tens of kW Server-level load buffering, BBU backup VII
Facility-level Traditional UPS (TUPS) Cycles (<<20 ms) 100s kW–MW Load protection, ride-through III
Grid-interactive UPS (GiUPS) 0.5–10 s (FFR) 100s kW–MW FFR, VRT, peak shaving, frequency reg. III
Grid-scale BTM BESS Minutes–hours MW–100s MW Peak shaving, RES integration, backup IV
FTM BESS Minutes–hours MW–GW Grid ancillary services, reserve IV
Backup generation Fuel Cell (FC) Seconds–minutes 100s kW–MW Low-carbon backup, co-generation V
Cooling support Thermal Energy Storage (TES) Minutes–hours MW thermal Cooling load shifting, RES support VI
Cost-optimized SLBESS Seconds–hours MW–100s MW BTM/FTM backup, cost-effective BESS VIII
@@ II-C Cooling Systems
TABLE II: Overview of GiUPS Operating Modes
@@ III-B1 Operational Modes of GiUPS
GiUPS can operate in several modes depending on the grid conditions and external control signals, including standard operation, discharge modes (full and partial disconnection), recharge mode, and energy export (bi-directional) mode [30, 34]. When a frequency variation is detected by an external controller, the GiUPS adjusts its operation to provide both positive and negative regulation by charging or discharging its batteries within operational limits. Table II, shows an overview of GiUPS operation modes.
@@ III-C BESS Sizing in GiUPS Systems
GiUPS BESS sizing must balance two objectives: maintaining backup reliability and delivering grid services such as FFR, primary regulation, and peak shaving [45, 106]. While backup needs establish the baseline energy capacity, grid-service participation adds important power-rating and cycling requirements, since services like frequency regulation are more power-intensive than energy-intensive. Economic sizing is therefore strongly influenced by cycle life, cumulative energy throughput, temperature, and average state of charge, because frequent shallow cycling can accelerate degradation [102, 57, 145, 132, 58, 95, 73, 32]. Moderate oversizing can reduce effective DoD and extend battery lifetime, but excessive oversizing increases capital cost, footprint, and thermal management burden without proportional benefit. Accordingly, optimal sizing should jointly consider backup energy requirements, committed grid-service power capacity, and degradation-aware lifecycle cost modeling. For AI DCs, this process must also account for additional reserved capacity for short-term load smoothing, since this function helps mitigate upstream ramp rates and transformer stress but should not compromise guaranteed backup autonomy.
@@ IV-A BESS Control Modes
TABLE III: Summary of power-smoothing architectures for AI data centers: benefits and challenges [105]
@@ IV-B2 Power Smoothing
As shown in Fig. 3, the GFM BESS effectively smooths the power fluctuations of the AI DCs resulting from training and inference jobs. An alternative approach is a hybrid E-STATCOM and BESS configuration, which combines the fast, high-speed charge/discharge response of supercapacitors with the higher energy capacity of BESS to achieve effective smoothing and sustained load support. This hybrid architecture can enhance resilience by improving demand flexibility, reducing flicker, and enabling broader grid-service capability. The primary trade-off is increased cost and footprint compared with standalone solutions. In addition, standalone supercapacitor-based E-STATCOM solutions can mitigate high-frequency AI load fluctuations and improve power quality for sensitive IT equipment by reducing flicker, although they provide limited long-duration energy support [105]. Table III compares the benefits and challenges of representative BESS-based and hybrid strategies for enabling grid support from AI DCs.